Set P1 =2.2 kN , P2_2 = 5 kN

Find the following and determine if they are in compression or tension

FAB, FAE, FBC, FBE,FCD, FCE, FDE
Answer #1

EF = 0 = Ax=0 AytOy = 2.2+5= 7.2kN Emo-o 0 4:4312KN(C) 2.319KN () 5.2705K Dy(12) = 2.2(0)+518) (T) 57512xN() 7.0667 KN - 3.13